Skip to end of metadata
Go to start of metadata

When wiki.ucar.edu was down I put misc entries here:

6-4-09 (twh) NOTE: I found (see below) that the 8m sonic swap on the profile tower appears to have occurred on 8-11 rather than 8-12. Oncley looked at photos of his jaunt to Moro Bay and found that he went there on 8-10. So the dates on the entries below should be changed to the previous day.

8/11 (twh: actually 8/10):
New heights were 5.37m above ref4 level (which is 0.04m above ref3 level).
Thus, the new heights should be 5.37+1.65=7.02m and 8.02m. upwind was setup
this way.
These are close enough to the existing profile 7.08 and 8.05, that we're not planning
to move them.

Yesterday (8/10), measured guy wire tensions on east horizontal array (with all the sonics now)
in a 6 m/s wind from the N:
SE 250
SW 220
NW 190
NE 240
These seem a bit low to us, but are reasonably consistent.

~12:00: Found that the interface board to 11b had pulled out of pressure2's
motherboard. Reset.

12:10-12:30:

  • Cycled power twice to upwind to get Pocketec running. Replaced the
    Pocketec in the process.
  • Swapped 4G sticks in pressure1 and pressure2
  • Have now downloaded all of these drives/sticks still need merge

~13:00 changed config. Restarted statsproc. covar_redo to go back to beginning
of this config. Config has new horiz array heights and new names for
some pressure sensors. Still need new heights for profile...

~13:30 We're off to Moro Bay for some R&R&R...

I sometimes get the following error when trying "aup", but a second "aup" always works:
Starting adam code.../usr/local/isff/scripts/adam/start_adam: line 30: 11238 Illegal instruction dsm '$ISFF/projects/$PROJECT/ISFF/config/'$CONFIG_XM
L
.
[Gordon says that this is an artifact of moving code from Viper RAM to SD card. If second
"aup" works, live with it.]

8/12 (twh: actually 8/11):

0945 - unpinched tubing to P.bedard
0950 - turned "profile" adam off to swap serial card
1005 - turned "profile" back on - don't have 8M card (rather than 8P)
1020-1030 - moved bedard port from on stake near ground to 13b boom (back a bit
from sonic)
1106 - added "adnaup" crontab on profile to restart every hour at 15 min after the hour.
~1130-1230:

  • measured all heights on profile:

1150-1210: replaced CSAT 720 with 733 at profile.8m (since high? speeds). (733 was the original "low" sonic at 8m when I arrived here, I think...)

6-4-09 (twh) I found a corresponding break in the 8m sonic data on 8/11, not 8/12

1210-1215: decided that it wasn't worth moving TRH's by 2cm in height. (7m wouldn't fit 2cm lower
anyway.)
1215: secured profile adam box